Complete the ordered pairs for the equation y = 8x + 8.

( ,0), (3/8, ), (0, ), (-3/8, )
No book. Thanks so much!!!


Answer by tutorcecilia(2144)

y = 8x + 8 [Plug-in the given values of either x or y]
(x, y) [Ordered pair]
.
Let y=0 since "0" represents the y value in point (x, 0)
0=8x+8 [Solve for x]
0-8=8x+8-8
-8=8x
-8/8=8x/8
-1=x
checking:
y = 8x + 8.
0=(8)(-1)+8
0=-8+8
0=0 [checks out]
.
.
Let x = 3/8 since it represents the x value in the point (3/8,y)
y = 8x + 8
y=(8)(3/8)+8 [Solve for y]
y=3+8
y=11
Plug-in ((3/8, 11) back into the original equatiion to check:
y = 8x + 8
11 = (8)(3/8)+8
11=3+8
11=11 [Checks out]
.
.

(0, y)
y = 8x + 8
y=(8)(0)+8
y=0+8
y=8
Checking:
y = 8x + 8.
8=(8)(0)+8
8=8 [checks out]
.
.
(-3/8,y)
y = 8x + 8
y = (8)(-3/8) + 8
y = -3+8
y=5
Checking:
5 = 8(-3/8) + 8
5 = -3+8
5=5 [Checks out]